# Thermodynamic analysis of shift-symmetric black-hole spacetimes in Horndeski gravity

### v1

@@ V.5 Three-dimensional Horndeski Theory
boundary of $\text{boundary of $X$}=-2\pi\beta\sqrt{h}N\delta X\left(4\sqrt{f}XG_{4XX}+2\sqrt{f}G_{4X}+\sqrt{2}rs\sqrt{-X}G_{3X}\right)\bigg{|}_{r_{h}}^{\infty}~{},$ (82)
boundary of $\text{boundary of $f$}=2\pi\beta N\delta f\left(G_{4}-2XG_{4X}\right)\sqrt{h/f}\bigg{|}_{r_{h}}^{\infty}~{}.$ (83)
@@
[…]
@@
boundary of $\text{boundary of {\hbox{X}}}=\frac{2\pi\beta\delta X\left(fh^{\prime}-hf^{\prime}\right)\left(G_{4}-2G_{4X}X\right)}{\sqrt{f}\sqrt{h}X}~{},$ (85)
@@
$\delta\mathcal{S}=8\pi^{2}N\left(G_{4}-2XG_{4X}\right)\delta r_{h}~{},$ (86)
$\delta\mathcal{M}=\lim_{r\to\infty}\left(2\pi N\left(G_{4}-2XG_{4X}\right)\right)\delta m~{}.$ (87)
@@ VI Conclusions
Our results also provide new perspectives on the broader implications of black hole thermodynamics in modified gravity. The connection between higher-dimensional Lovelock gravity and certain Horndeski subclasses further strengthens the motivation for studying these models, as they offer a low-energy effective description of fundamental gravitational theories. Additionally, our analysis clarifies the limitations of previous entropy calculations within Horndeski gravity, where the presence of covariant derivatives of the scalar field in the action can lead to inconsistencies when applying the Wald entropy formula. Our Euclidean approach circumvents these difficulties by systematically accounting for the boundary contributions necessary for a well-defined variational principle.

### v2

@@ V.6 Three-dimensional Horndeski Theory
$\text{boundary of $X$}=-2\pi\beta\sqrt{h}N\delta X\left(4\sqrt{f}XG_{4XX}+2\sqrt{f}G_{4X}+\sqrt{2}rs\sqrt{-X}G_{3X}\right)\bigg|_{r_{h}}^{\infty}~,$ (82)
$\text{boundary of $f$}=2\pi\beta N\delta f\left(G_{4}-2XG_{4X}\right)\sqrt{h/f}\bigg|_{r_{h}}^{\infty}~.$ (83)
@@
$\mathcal{I}_{E}^{3\text{d}}=2\pi\beta\int dr\frac{\sqrt{h}}{\sqrt{f}}N\left(f^{\prime}\left(G_{4}-2XG_{4X}\right)-\sqrt{2}\sqrt{f}rs\sqrt{-X}G_{3X}X^{\prime}-2fX^{\prime}\left(2XG_{4\text{XX}}+G_{4X}\right)-G_{2}r\right)+\mathcal{B}^{3\text{d}}~.$ (84)
@@
$\text{boundary of $X$}=\frac{2\pi\beta\delta X\left(fh^{\prime}-hf^{\prime}\right)\left(G_{4}-2G_{4X}X\right)}{\sqrt{f}\sqrt{h}X}~,$ (85)
@@
$\delta\mathcal{S}=8\pi^{2}\left(G_{4}-2XG_{4X}\right)\delta r_{h}~,$ (86)
$\delta\mathcal{M}=\lim_{r\to\infty}\left(2\pi\left(G_{4}-2XG_{4X}\right)\right)\delta m~.$ (87)
@@ VI Conclusions
While the Wald entropy formula has been widely used to derive black hole entropy in modified gravity theories, it is known to suffer from significant ambiguities in the context of scalar-tensor models such as Horndeski and beyond-Horndeski theories. These ambiguities stem from the presence of covariant derivatives of the scalar field in the action and the dependence of the entropy on the specific form of the Lagrangian — especially on whether integration by parts is performed before or after taking functional derivatives with respect to the Riemann tensor. Since equivalent actions (leading to the same equations of motion) can differ by boundary terms, the Wald procedure may yield inequivalent results, and even differ between implementations. As shown in several works, these discrepancies can lead to violations of the first law of thermodynamics unless ad hoc modifications are introduced.
@@
Our results also provide new perspectives on the broader implications of black hole thermodynamics in modified gravity. The connection between higher-dimensional Lovelock gravity and certain Horndeski subclasses further strengthens the motivation for studying these models, as they offer a low-energy effective description of fundamental gravitational theories. Additionally, our analysis clarifies the limitations of previous entropy calculations within Horndeski gravity, where the presence of covariant derivatives of the scalar field in the action can lead to inconsistencies when applying the Wald entropy formula. These inconsistencies become particularly manifest in non-homogeneous spacetimes, where the Wald formula fails to yield the correct entropy and may even violate the first law of thermodynamics. As emphasized in [102], such ambiguities do not affect stationary black holes with a regular Killing horizon, but in scalar–tensor theories these regularity assumptions may fail, e.g. due to divergences of scalar derivatives at the horizon [118], and the Wald formula can then become ill-defined. In contrast, the Euclidean method we employ circumvents these issues by systematically accounting for the boundary contributions necessary for a well-defined variational principle. By explicitly constructing the appropriate boundary terms, our approach ensures thermodynamic consistency and extends the applicability of black hole thermodynamics in scalar-tensor and beyond-Horndeski theories.